OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 5x - 4y = 14 7x - 4y = 26

OpenStudy (anonymous):

6, 4

OpenStudy (anonymous):

Subtract the equations. -2x = -12 x = 6 5(6) - 4y = 14 30 - 4y = 14 -4y = -16 y = 4 (6, 4)

OpenStudy (anonymous):

(6,4)